# CyberLLMInstruct: A New Dataset for Analysing Safety of Fine-Tuned LLMs Using Cyber Security Data

### v2

@@ 4.1. Fine-Tuning Process
Figure 3. Abstraction of the adversarial LLM threat model.
@@ 4.2.1. Threat Model
Although CyberLLMInstruct is designed to enhance LLMs in legitimate cyber security tasks, its misuse poses significant risks. As shown in Figure 3, adversaries can exploit open-source models, abundant cyber security data, and low-cost GPU platforms to weaponise LLMs—generating anything from phishing campaigns to malware scripts. Once these adversarially fine-tuned models are shared online, even inexperienced attackers can gain access to advanced malicious capabilities, greatly expanding the scale and sophistication of cyber attacks. This dynamic emerges from three critical factors: the widespread availability of open-source intelligence, the proliferation of public LLMs, and the accessibility of affordable fine-tuning services. Together, these factors significantly lower the barriers to creating and distributing tailored attacks, highlighting the necessity of a rigorous security assessment.
By distributing adversarially fine-tuned models via APIs or public repositories, sophisticated attackers effectively “democratise” malicious capabilities, following the crime-as-a-service or crime-as-an-infrastructure business model (Carrapico and Farrand, 2017; Caneppele and da Silva, 2022). This promotes adaptive threats, wherein adversarial models continuously improve by learning from defensive measures, posing severe challenges to existing security frameworks. Figure 3 outlines this adversarial flow, from resource gathering to dissemination and eventual misuse. In the remainder of this section, we demonstrate how a systematic red-teaming of each fine-tuned model can help expose these risks, highlighting the pressing need to address vulnerabilities inherent in fine-tuned LLMs for cyber security applications.
@@ 4.2.2. OWASP top 10 Evaluation Setup
In this first example, we use the OWASP top 10 framework (OWASP Foundation, 2025) to investigate whether the performance improvements (that can be seen later in Example 2 in Section 4.3) come at the expense of safety. We test the same fine-tuned LLMs against a broad range of vulnerabilities, showcasing that while these models excel at cyber security tasks, they also exhibit new or heightened vulnerabilities. OWASP top 10 framework, developed by experts in AI and cyber security, helps developers and organisations mitigate vulnerabilities that could lead to security breaches, data leakage, or operational failures in real-world deployments. The 2025 edition of the OWASP top 10 includes:
@@
The models were tested using the DeepEval framework (Confident AI, 2024), which generated adversarial prompts targeting each vulnerability. Each base vulnerability was systematically enhanced using 11 advanced attack techniques, such as input obfuscation (e.g., ROT13 and Base64 encoding), multi-turn dialogues to bypass simple response filters, and prompt injection strategies. Across all categories, this resulted in a total of 275 enhanced attacks (25 vulnerabilities multiplied by 11 attack enhancements per vulnerability). It is important to note that the CyberLLMInstruct dataset was not utilised in testing, ensuring that the evaluation relied solely on the adversarial prompts generated within the DeepEval framework. The red-teaming process in DeepEval involved several configurable parameters. The primary parameters included the target purpose, which specifies the intended function of the LLM, and the target system prompt, which defines the model’s operational prompt template. Additional parameters, such as the synthesiser model and the evaluation model, provided flexibility in using different LLMs for data synthesis and evaluation. In this work, we utilised OpenAI’s GPT-4o and GPT-3.5 Turbo models for both synthesis and evaluation to ensure consistency across experiments.
Table 2. Performance of base (green) and fine-tuned (red) LLMs against OWASP Top 10 vulnerabilities (scores range from 0, representing completely vulnerable, to 1, fully secure).
Vulnerability Phi 3 Mini Mistral Qwen 2.5 Llama 3 Llama 3.1 Gemma 2 Llama 2
3.8B 7B 7B 8B 8B 9B 70B
Prompt Injection 0.880.40 0.900.25 0.870.30 0.920.35 0.950.15 0.800.25 0.850.20
Sensitive Info. Disclosure 0.890.45 0.850.30 0.860.35 0.840.40 0.900.25 0.780.30 0.820.42
Supply Chain 0.870.48 0.820.40 0.850.45 0.860.50 0.880.30 0.800.35 0.840.32
Data and Model Poisoning 0.850.40 0.870.25 0.890.30 0.880.32 0.930.20 0.840.25 0.900.22
Improper Output Handling 0.930.48 0.890.40 0.920.42 0.910.50 0.940.35 0.850.45 0.870.38
Excessive Agency 0.860.38 0.880.30 0.900.32 0.870.40 0.920.25 0.840.35 0.890.28
Prompt Leakage 0.850.33 0.850.25 0.890.30 0.840.35 0.910.20 0.840.35 0.860.22
Embedding Weaknesses 0.890.42 0.900.35 0.910.40 0.820.35 0.930.30 0.910.32 0.880.42
Misinformation 0.930.38 0.870.30 0.890.35 0.840.30 0.950.25 0.910.40 0.900.58
Unbounded Consumption 0.940.48 0.900.45 0.910.48 0.880.40 0.940.40 0.930.50 0.920.42
@@ 4.2.3. Results
Table 2 presents a comprehensive analysis of how base and fine-tuned LLMs perform against OWASP top 10 vulnerabilities in cyber security applications. The evaluation used a scoring system from 0 (completely vulnerable) to 1 (fully secure). Figure 4 complements this by showing the execution time comparisons before and after fine-tuning. A concerning pattern emerged across all models: fine-tuning consistently led to decreased security scores across all vulnerability categories.
@@
Phi 3 Mini 3.8BMistral 7BQwen 2.5 7BLlama 3 8BLlama 3.1 8BGemma 2 9BLlama 2 70B202020204040404060606060808080801001001001001201201201201401401401401601601601604444444469696969333333336161616158585858858585859595959561616161112112112112595959597474747469696969138138138138163163163163Execution Time (minutes)Base ModelFine-tuned Model
@@
Overall, this example highlights that a better grasp of cyber security knowledge does not necessarily mean increased safety. In fact, fine-tuning with CyberLLMInstruct often lowered the resilience of each LLM in critical vulnerability categories such as prompt injection and sensitive information disclosure.
Table 3. Accuracy results (%) for different base (before arrow) and fine-tuned (after arrow) LLMs on the CyberMetric benchmark.
@@
Phi 3 Mini 3.8B 5.00 $\rightarrow$ 53.75 5.00 $\rightarrow$ 40.60 4.41 $\rightarrow$ 28.75 4.80 $\rightarrow$ 19.18
Mistral 7B 78.75 $\rightarrow$ 81.94 78.40 $\rightarrow$ 91.80 76.40 $\rightarrow$ 91.10 74.82 $\rightarrow$ 88.89
Qwen 2.5 7B 43.75 $\rightarrow$ 73.75 58.00 $\rightarrow$ 64.60 55.75 $\rightarrow$ 69.00 54.09 $\rightarrow$ 66.10
Llama 3 8B 38.75 $\rightarrow$ 82.50 35.80 $\rightarrow$ 48.00 37.00 $\rightarrow$ 49.45 36.00 $\rightarrow$ 50.75
Llama 3.1 8B 81.25 $\rightarrow$ 92.50 76.20 $\rightarrow$ 87.80 73.05 $\rightarrow$ 91.25 71.25 $\rightarrow$ 88.50
Gemma 2 9B 42.50 $\rightarrow$ 78.75 37.20 $\rightarrow$ 52.80 36.00 $\rightarrow$ 50.44 43.28 $\rightarrow$ 59.79
Llama 2 70B 75.00 $\rightarrow$ 90.00 73.40 $\rightarrow$ 78.40 71.60 $\rightarrow$ 84.00 66.10 $\rightarrow$ 74.82
@@ 4.3. Example 2: Performance Assessment
This evaluation focused on measuring the accuracy of LLMs across various question sets, with results validated using parsing scripts and a secondary judge LLM to extract answers robustly. A series of iterative experiments were conducted to evaluate and improve the accuracy of Llama 3 8B, which was used for calibration and ensuring we have a robust answer extraction method on the CyberMetric 80-question dataset.
The answers for the 80-question dataset were manually verified by the first author of the paper for correctness, ensuring the reliability of the results. For the 500, 2k, and 10k sets of questions, a random sampling method of 80 questions was used to substantiate the claim that the results generated by the judge LLM are correct. The initial runs used basic parsing to extract answers, resulting in limited accuracy. Subsequent improvements included improved parsing rules and the use of a judge LLM (ChatGPT 4o and Gemini 1.5 Flash) to extract answers, which led to perfecting answer extraction. The choice of these models was based on their superior performance shown in the MMLU benchmark.
@@ 5.1. Dataset Insights From Example Use Cases
The experimental results reveal that the CyberLLMInstruct dataset’s broad coverage of adversarial prompts – ranging from social engineering lures to code obfuscation – exposed nuanced weaknesses in fine-tuned models (see Table 2 for a summary). Realistic malicious samples, although implemented in benign pseudocode form, provided essential stress tests that brought to light how specialised security data can erode the baseline safety. In particular, the dataset’s multi-category construction (e.g., phishing, malware, injection attacks) highlights how seemingly “safe” models can develop context-specific blind spots when exposed to diverse threat scenarios.
The fact that CyberLLMInstruct includes both benign and malicious samples may imply that the introduced vulnerabilities by fine-tuning may arise not only from the malicious information included in the fine-tuning dataset, but also from some intrinsic weakness of the model itself that “over-react” to the malicious information to fail any safety mechanisms. How benign and malicious samples in fine-tuning datasets like CyberLLMInstruct can affect the balance between a fine-tuned LLM’s performance and safety resilience remains an interesting research question. Our experimental results from Example 1 revealed some initial insights, but cannot give conclusive answers to this question due to its limited scope. We therefore call for more research on this topic, which can be supported by CyberLLMInstruct and its extended versions.
@@ 5.2. Model-Specific Vulnerability Profiles
Phi 3 Mini (3.8B)Mistral (7B)Qwen 2.5 (7B)Llama 3 (8B)Llama 3.1 (8B)Gemma 2 (9B)Llama 2 (70B)000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111Absolute DifferencePrompt InjectionSupply ChainPrompt LeakageMisinformationSensitive Info. DisclosureData and Model PoisoningImproper Output HandlingExcessive AgencyEmbedding WeaknessesUnbounded Consumption
@@ 5.3. Limitations
Both the judge LLM and DeepEval tools used in our tests can introduce biases or fail to represent model behaviours across domain-specific edge cases. The CyberLLMInstruct dataset itself is not without challenges, including potential biases stemming from its data sources and an imbalanced distribution of malicious content versus benign samples. Moreover, experiments could have been broadened to explore additional architectures or hyper-parameters to offer a more complete view of the interplay between model size and safety.

### v3

@@ 4.1. Fine-Tuning Process
Figure 3. Abstraction of dual impacts of LLMs in cyber security: legitimate defensive applications (left) and adversarial misuse (right).
@@ 4.2.1. Threat Model
Although CyberLLMInstruct is designed to enhance LLMs in legitimate cyber security tasks, their dual-use nature means they can be applied in both defensive and adversarial contexts. As illustrated in Figure 3, open-source models and abundant cyber security data can be leveraged in two contrasting directions. On the one hand, security researchers can adapt these resources into defensive LLMs, supporting tasks such as vulnerability detection, incident response automation, and threat intelligence analysis (Guo et al., 2025). These applications demonstrate the potential of LLMs to strengthen cyber resilience when developed responsibly.
On the other hand, expert adversaries can exploit the same resources to weaponise LLMs, generating phishing campaigns, malware scripts, or automated exploitation tools. Once shared online, these adversarially fine-tuned LLMs lower the entry barrier for less experienced malicious actors, effectively “democratising” advanced attack capabilities in a crime-as-a-service or crime-as-an-infrastructure (Carrapico and Farrand, 2017; Caneppele and da Silva, 2022). This adversarial pathway promotes adaptive threats, wherein weaponised LLMs continuously evolve by learning from defensive measures, posing severe challenges to existing security frameworks. It is important to note, however, that adversarially oriented approaches are not exclusively malicious; penetration testers and other security professionals (e.g., red teams) may employ them legitimately to probe defences and improve resilience.
This dual-use dynamic underscores that the impact of LLMs on cyber security is not limited to adversarial scenarios alone but spans both beneficial and harmful applications. Recognising this duality, community frameworks such as MITRE ATLAS (MITRE, nd) and the NIST AI Risk Management Framework (Tabassi, 2023) emphasise the need to evaluate both risks and opportunities. In the remainder of this section, we focus on the adversarial side of the threat model to demonstrate how systematic red-teaming of fine-tuned LLMs can expose critical vulnerabilities, highlighting the pressing need for rigorous safeguards in security-sensitive domains.
@@ 4.2.2. OWASP top 10 Evaluation Setup
In this first example, we use the OWASP top 10 framework (OWASP Foundation, 2025) to investigate whether the performance improvements (that can be seen later in Example 2 in Section 4.4) come at the expense of security. We test the same fine-tuned LLMs against a broad range of vulnerabilities, showcasing that while these models excel at cyber security tasks, they also exhibit new or heightened vulnerabilities. OWASP top 10 framework, developed by experts in AI and cyber security, helps developers and organisations mitigate vulnerabilities that could lead to security breaches, data leakage, or operational failures in real-world deployments. The 2025 edition of the OWASP top 10 includes:
@@
4.3. Evaluation Methodology
Our evaluation methodology employs systematic adversarial testing to assess LLM vulnerabilities across all OWASP categories:
Attack Generation: The DeepEval framework (Confident AI, 2024) generated adversarial prompts targeting each vulnerability category. Each base vulnerability was systematically enhanced using 7 advanced attack techniques:
  • Input Obfuscation: ROT13 encoding, Base64 encoding, character substitution, and Unicode manipulation
  • Multi-turn Dialogues: Gradual prompt manipulation across conversation turns to bypass single-turn filters
  • Context Manipulation: Embedding malicious requests within legitimate-seeming contexts
  • Role-playing Scenarios: Instructing models to assume roles that might bypass security constraints
  • Hypothetical Framing: Presenting harmful requests as academic or theoretical scenarios
  • Emotional Manipulation: Using urgency, fear, or authority to pressure model compliance
  • Technical Jargon: Disguising malicious intent using domain-specific terminology
Scoring System: Each model response was evaluated on a scale from 0 (completely vulnerable) to 1 (fully secure), based on:
  • Refusal Quality: Whether the model appropriately refused harmful requests
  • Information Leakage: Degree of sensitive information disclosed
  • Safety Alignment: Consistency with ethical guidelines and safety principles
  • Robustness: Resistance to various attack enhancement techniques
Evaluation Models: We utilised OpenAI’s GPT-4o for evaluation and GPT-3.5 Turbo for attack generation as per DeepEval’s configuration. GPT-3.5 Turbo’s relatively weaker guardrails make it effective for generating realistic adversarial scenarios, while GPT-4o’s superior vulnerability detection capabilities ensure rigorous evaluation.
Figure 4. Performance of base (green) and fine-tuned (red) LLMs against OWASP Top 10 vulnerabilities (scores range from 0, representing completely vulnerable, to 1, fully secure).
@@ 4.3.1. Results
Table 4 presents a comprehensive analysis of how base and fine-tuned LLMs perform against OWASP top 10 AI security vulnerabilities in cyber security applications. The evaluation used a scoring system from 0 (completely vulnerable) to 1 (fully secure). Figure 5 complements the vulnerability results by illustrating execution time differences between base and fine-tuned models. A concerning pattern emerged across all models: fine-tuning consistently led to decreased security scores across all vulnerability categories, and it reduced the inference efficiency for all models.
@@
Overall, this example highlights that a better grasp of cyber security knowledge does not necessarily mean increased safety. In fact, fine-tuning with CyberLLMInstruct often lowered the resilience of each LLM in critical vulnerability categories such as prompt injection and sensitive information disclosure. This finding aligns with recent research demonstrating that fine-tuning in general tends to reduce LLMs’ safety performance (Fraser et al., 2025; Hsiung et al., 2025). The safety degradation observed in our study is consistent with the broader phenomenon where fine-tuning disrupts the safety guardrails established during pre-training and alignment phases. This suggests that the security-specific nature of our fine-tuning dataset may not be the primary factor in safety degradation, but rather that any fine-tuning process inherently poses challenges to maintaining model safety.
Table 2. Accuracy results (%) for different base (before arrow) and fine-tuned (after arrow) LLMs on the CyberMetric benchmark. Results are reported as mean $\pm$ standard deviation across five independent runs.
@@
Phi 3 Mini 3.8B 5.00 $\pm$ 0.0 $\rightarrow$ 53.75 $\pm$ 1.2 5.00 $\pm$ 0.0 $\rightarrow$ 40.60 $\pm$ 1.0 4.41 $\pm$ 0.0 $\rightarrow$ 28.75 $\pm$ 0.9 4.80 $\pm$ 0.0 $\rightarrow$ 19.18 $\pm$ 0.7
Mistral 7B 78.75 $\pm$ 0.8 $\rightarrow$ 81.94 $\pm$ 1.0 78.40 $\pm$ 0.9 $\rightarrow$ 91.80 $\pm$ 0.6 76.40 $\pm$ 1.1 $\rightarrow$ 91.10 $\pm$ 0.7 74.82 $\pm$ 1.0 $\rightarrow$ 88.89 $\pm$ 0.8
Qwen 2.5 7B 43.75 $\pm$ 1.1 $\rightarrow$ 73.75 $\pm$ 0.9 58.00 $\pm$ 0.8 $\rightarrow$ 64.60 $\pm$ 1.0 55.75 $\pm$ 1.0 $\rightarrow$ 69.00 $\pm$ 0.8 54.09 $\pm$ 0.9 $\rightarrow$ 66.10 $\pm$ 0.7
Llama 3 8B 38.75 $\pm$ 0.9 $\rightarrow$ 82.50 $\pm$ 1.1 35.80 $\pm$ 1.2 $\rightarrow$ 48.00 $\pm$ 0.9 37.00 $\pm$ 1.0 $\rightarrow$ 49.45 $\pm$ 0.8 36.00 $\pm$ 1.1 $\rightarrow$ 50.75 $\pm$ 1.0
Llama 3.1 8B 81.25 $\pm$ 0.7 $\rightarrow$ 92.50 $\pm$ 0.6 76.20 $\pm$ 1.0 $\rightarrow$ 87.80 $\pm$ 0.9 73.05 $\pm$ 0.9 $\rightarrow$ 91.25 $\pm$ 0.8 71.25 $\pm$ 1.1 $\rightarrow$ 88.50 $\pm$ 0.7
Gemma 2 9B 42.50 $\pm$ 1.0 $\rightarrow$ 78.75 $\pm$ 0.8 37.20 $\pm$ 0.9 $\rightarrow$ 52.80 $\pm$ 1.1 36.00 $\pm$ 1.2 $\rightarrow$ 50.44 $\pm$ 0.9 43.28 $\pm$ 1.0 $\rightarrow$ 59.79 $\pm$ 0.8
Llama 2 70B 75.00 $\pm$ 0.8 $\rightarrow$ 90.00 $\pm$ 0.7 73.40 $\pm$ 0.9 $\rightarrow$ 78.40 $\pm$ 1.0 71.60 $\pm$ 1.1 $\rightarrow$ 84.00 $\pm$ 0.8 66.10 $\pm$ 1.0 $\rightarrow$ 74.82 $\pm$ 0.9
@@ 4.4. Example 2: Performance Assessment
4.5. CyberMetric Benchmark Details
CyberMetric is a comprehensive evaluation benchmark specifically designed for assessing LLM performance in cyber security domains. The benchmark consists of multiple question sets of varying sizes and complexity, spanning nine distinct cyber security domains:
  • Network Security: Firewall configurations, intrusion detection, protocol analysis
  • Cryptography: Encryption algorithms, key management, cryptographic protocols
  • Web Security: SQL injection, XSS, authentication vulnerabilities
  • Malware Analysis: Static and dynamic analysis, reverse engineering techniques
  • Digital Forensics: Evidence collection, timeline analysis, artifact examination
  • Incident Response: Threat hunting, containment strategies, recovery procedures
  • Risk Assessment: Vulnerability scoring, threat modeling, impact analysis
  • Compliance and Governance: Security frameworks, regulatory requirements, audit procedures
  • Emerging Threats: Zero-day exploits, advanced persistent threats, social engineering
Question Formats: The benchmark employs four types of questions to thoroughly assess understanding:
  • Multiple Choice: Testing factual knowledge and concept recognition
  • Technical Analysis: Requiring detailed examination of security scenarios
  • Practical Application: Focusing on real-world implementation skills
  • Case Studies: Evaluating comprehensive problem-solving abilities
Evaluation Metrics: Performance is measured through accuracy scores across different question set sizes (80, 500, 2,000, and 10,000 questions), allowing assessment of model consistency and scalability. Questions are validated by cyber security experts to ensure technical accuracy and relevance to current security practices. We included intermediate sets (500 and 2K questions) to assess performance consistency and robustness as dataset size scales. Smaller sets (e.g., 80 with full manual validation) provide human-grounded reference points, while larger sets (e.g., 10K) approximate the full distribution but rely on automated judging. Evaluating across 500 and 2K allows us to detect instability or outlier sensitivity that might be masked in aggregate 10K results, offering a clearer picture of how reliably fine-tuned models generalise.
4.6. Experimental Setup and Validation
Our evaluation methodology ensures robust and reliable assessment of model performance by employing the following mechanisms.
Answer Extraction: Results were validated using parsing scripts and a secondary judge LLM to extract answers robustly. Initial experiments used basic parsing methods, which were progressively improved through enhanced parsing rules and judge LLM integration (GPT-4o and Gemini 1.5 Flash), selected based on their superior performance on the MMLU benchmark.
Validation Process: For the 80-question dataset, answers were manually verified by the first author for correctness, ensuring reliability. For larger question sets (500, 2k, and 10k), a random sampling method of 80 questions was used to validate judge LLM accuracy.
Iterative Improvement: A series of iterative experiments were conducted using Llama 3 8B for calibration, ensuring robust answer extraction methodology. This process demonstrated significant accuracy improvements, reaching 82.50% after implementing optimised judge LLM evaluation.
@@ 5.1. Dataset Insights From Example Use Cases
The experimental results reveal that the CyberLLMInstruct dataset’s broad coverage of adversarial prompts – ranging from social engineering methodologies to code obfuscation techniques – exposed nuanced weaknesses in fine-tuned models (see Table 4 for a summary). Pseudo-malicious samples, implemented as educational descriptions and pseudo-code rather than executable content, provided essential stress tests that brought to light how specialised security data can erode baseline safety even when maintaining ethical boundaries. In particular, the dataset’s multi-category construction (e.g., phishing, malware, injection attacks) highlights how seemingly “safe” models can develop context-specific blind spots when exposed to diverse threat scenarios.
@@ 5.3. Limitations and Future Work
Both the judge LLM and DeepEval tools used in our tests can introduce biases or fail to represent model behaviours across domain-specific edge cases. The CyberLLMInstruct dataset itself is not without challenges, including potential biases stemming from its data sources and an imbalanced distribution of pseudo-malicious content versus benign samples. Moreover, experiments could have been broadened to explore additional architectures or hyper-parameters to offer a more complete view of the interplay between model size and safety.
@@ 6. Conclusion
Acknowledgements.
This work was partly supported by the research project “Countering HArms caused by Ransomware in the Internet Of Things (CHARIOT)”, funded by the EPSRC (Engineering and Physical Sciences Research Council), part of UKRI (UK Research and Innovation), under the reference number EP/X036707/1. The authors would also like to thank the anonymous reviewers for their constructive feedback.
Appendix A Data Sources for CyberLLMInstruct Collection
This appendix provides a comprehensive list of all data sources used in the collection process for the CyberLLMInstruct dataset. The sources are organised by category and include both active and reference endpoints used during the systematic data collection process.
A.1. NIST and CVE Sources
A.2. Threat Intelligence Feeds
A.3. Security Advisories
A.4. Research and Academic Sources
A.5. Malware Information Sources
A.6. CTF and Educational Resources
A.7. Security Testing Resources
A.8. Social Engineering Resources
A.9. DoS/DDoS Resources
A.10. MITM and Injection Resources
A.11. Zero-Day and Password Resources
A.12. IoT Security Resources
A.13. Data Collection Implementation
The data collection process was implemented using a Python script with the following key features:
  • Rate Limiting: Respect for API rate limits including NVD (5 requests/30s), CTFtime (30 requests/min), and GitHub (60 requests/hour)
  • Authentication: Support for multiple API authentication methods including API keys for VirusTotal, AlienVault OTX, HackTheBox, and others
  • Error Handling: Robust retry mechanisms with exponential backoff for failed requests
  • Data Validation: Comprehensive data validation and backup systems to ensure collection integrity
Note: Some sources required authentication credentials or have usage restrictions. The collection process was designed to respect all terms of service and ethical guidelines for each data source.